The perimeter of a rectangle is 44ft. The length of the rectangle is 2ft less than three times the width. Find the length of the rectangle.
Answer by iheartcrmnljstce(1)

The perimeter is 44 ft.
Perimeter is the two lengths and two widths added together.
Let each width be x.
Each length is 2ft less than 3 times the width...
...2ft less than ...means you are subtracting from the 3 times the width
3 times the width equals 3x
then you subtract 2ft...
Each length is 3x - 2
So you must add two widths and two lengths to equal the perimeter...
x+x+(3x-2)+(3x-2)=44
add similiar variables together...
x+x+3x+3x=8x
-2-2= -4
which equals... 8x-4=44
+4 +4
8x=48
divide each side by 8
x=6
then you need to plug in x to both the widths and lengths...
each width is 6
each length is (3(6)-2) --->18-2=16

check answer by adding them together... 6+6=12, 16+16=32... 12+32=44
*
*
*


The answer is each length is 16.